What is the mass of 3.25 moles of Cu(OH)2
1
Expert's answer
2013-12-21T11:24:59-0500
The mass of compound can be obtained from the equation:
m = n*M
where n is number of moles, M is a molar mass of compound.
M[Cu(OH)2] = 63.546+(15.999+1.008)*2 = 97.56 g/mol
m[Cu(OH)2] = 3.25*97.56 = 317 g